OpenStudy (anonymous):

Find the slope, if possible, of the line passing through each pair of points. (2,-1) and (6, 3)

OpenStudy (anonymous):

It passes through but I do not understand, can someone explain?

hero (hero):

There are a few ways of finding the slope: 1. Directly (using slope formula) 2. Graphically (using rise/run) 3. Algebraically (using y = mx + b)

OpenStudy (anonymous):

algebraically please if you don't mind

hero (hero):

If you want to find it algebraically, you have to insert each point, one at a time, into the equation y = mx + b. Doing that, you'll get -1 = 2m + b 3 = 6m + b

OpenStudy (anonymous):

Oops I am sorry I have to use the slope formula.

hero (hero):

Okay, then use this slope formula: \[m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

slope = (y2 - y1)/(x2 - x1)?

OpenStudy (anonymous):

slope = (3 - (-1))/(6 - 2) slope = (3 + 1)/4 slope = 4/4 slope = 1

OpenStudy (anonymous):

I think this is right

hero (hero):

Good job. Now, here's a challenge. See if you can figure it out the algebraic way

OpenStudy (anonymous):

good golly I want to learn to like algebra

OpenStudy (anonymous):

I will try but I have more homework to do.